OpenStudy (anonymous):

Is the final answer to x^2 = -5... x=i5 and x=-i5? Or am I missing something here?

OpenStudy (anonymous):

root 5 also -5 = -1 * 5 take root i root5

OpenStudy (anonymous):

\[i \sqrt{5}\]

OpenStudy (anonymous):

i radical 5 and - i radical 5, right?

OpenStudy (anonymous):

just i not - i \[\sqrt{5} \times i\]

OpenStudy (anonymous):

\[i = \sqrt{-1}\]

OpenStudy (anonymous):

Oh ok. Thanks!

OpenStudy (anonymous):

you are welcome